OpenStudy (anonymous):

The time cake a must bake is between 25 and 30 minutes, inclusive.

OpenStudy (anonymous):

@wio could you please help me?

OpenStudy (anonymous):

What is the question?

OpenStudy (anonymous):

I'm dealing with compound inequalities, that is the question

OpenStudy (anonymous):

Inclusive means use \(\leq \)

OpenStudy (anonymous):

\[ 25\leq t \leq 30 \]

OpenStudy (anonymous):

Ah, I was confused as to what inclusive meant.

OpenStudy (anonymous):

Would you mind helping me with another?

OpenStudy (anonymous):

Solve. 5 < x - 2 < 11

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

Add \(2\) to all sides.

OpenStudy (anonymous):

Okay, add 2 to 2,5, and 11?

OpenStudy (anonymous):

the question is to how to bake a cake @wio

OpenStudy (anonymous):

No, the question was how would I write that as a compound inequality

OpenStudy (anonymous):

@genius12

OpenStudy (anonymous):

Basically the 't', i.e. time to bake a cake is between 25 and 30 minutes inclusive. You write that the way @wio did in the previous post. @fenkjwfn